Chapman, Inc.'s Mexican subsidiary, V. Gomez Corporation, is expected to pay to Chapman 20 pesos in dividends in 1 year after all foreign and U.S. taxes have been subtracted. The exchange rate in 1 year is expected to be $0.08 per peso. After this, the peso is expected to depreciate against the dollar at a rate of 4% a year forever due to the different inflation rates in the United States and Mexico. The peso-denominated dividend is expected to grow at a rate of 5% a year indefinitely. Chapman owns 15 million shares of V. Gomez.

What is the present value of the dividend stream, in dollars, assuming V. Gomez's cost of equity is 14%? Do not round intermediate calculations. Round your answer to the nearest dollar.

$  
Answer #1

Dividend in year 1 in USD, D1 = 20 x 0.08 = $1.60

Growth in USD, g = (1 + 5%) x (1 - 4%) - 1 = 0.80%

PV of dividends = D1 / (r - g)

= 1.60 / (14% - 0.80%)

= $12.12 per share

Total Value = 12.12 x 15m = $181,818,182